# Requirements Coverage-Guided Minimization for Natural Language Test Cases

### v1

@@ Abstract.
As software systems evolve, test suites tend to grow in size and often contain redundant test cases. Such redundancy increases testing effort, time, and cost. Test suite minimization (TSM) aims to eliminate such redundancy while preserving key properties such as requirement coverage and fault detection capability. In this paper, we propose RTM (Requirement coverage-guided Test suite Minimization), a novel TSM approach designed for requirement-based testing (validation), which can effectively reduce test suite redundancy while ensuring full requirement coverage and a high fault detection rate (FDR) under a fixed minimization budget. Based on common practice in critical systems where functional safety is important, we assume test cases are specified in natural language and traced to requirements before being implemented. RTM preprocesses test cases using three different preprocessing methods, and then converts them into vector representations using seven text embedding techniques. Similarity values between vectors are computed utilizing three distance functions. A Genetic Algorithm, whose population is initialized by coverage-preserving initialization strategies, is then employed to identify an optimized subset containing diverse test cases matching the set budget.
We evaluate RTM on an industrial automotive system dataset comprising $736$ system test cases and $54$ requirements. Experimental results show that RTM consistently outperforms baseline techniques in terms of FDR across different minimization budgets while maintaining full requirement coverage. Furthermore, we investigate the impact of test suite redundancy levels on the effectiveness of TSM, providing new insights into optimizing requirement-based test suites under practical constraints.
@@ 1. Introduction
Software testing is a critical activity throughout the software development life cycle, playing a key role in early defect detection and compliance with requirements (Myers et al., 2011; Everett and McLeod Jr, 2007). In particular, testing driven by requirements is conducted to validate systems before their deployment (Dustin et al., 2009). This is especially critical for systems that must demonstrate functional safety, such as automotive systems. In such contexts, it is also common to specify test cases in natural language (following templates) to facilitate their analysis and traceability to requirements, before manually writing the corresponding test code.
@@
However, due to the increasing complexity of systems and their continuous change by large teams, test suites are hard to maintain, tend to grow, and are often redundant (i.e., some of the test cases are likely to detect the same faults), thus leading to waste of time and resources (Rothermel et al., 2002). Further, in practice, testers have limited resources and time to test new system updates. To address this, given a test budget, test suite minimization (TSM) is proposed to prune test cases that are most likely to be redundant, while satisfying coverage criteria and maximizing fault detection (Harrold et al., 1993; von Ronne, 1999), thus significantly reducing testing cost and time while limiting the impact on system quality or compliance.
@@
While a number of approaches have been proposed for TSM (Noemmer and Haas, 2019; Miranda and Bertolino, 2017; Coviello et al., 2018; Liu et al., 2011; Viggiato et al., 2022; Cruciani et al., 2019; Zhang et al., 2019; Hemmati et al., 2013; Pan et al., 2023, 2024; Marchetto et al., 2017; Xia et al., 2021; Anwar et al., 2019; Yoo and Harman, 2010), most existing techniques rely on code and structural coverage criteria (e.g., statement or branch coverage). These techniques do not address the unique challenges of minimizing test suites driven by system requirements, with test cases often specified in natural language. Moreover, to the best of our knowledge, no existing solution is capable of simultaneously maintaining full requirement coverage and adhering to a fixed minimization budget, a common situation in industrial practice due to limited time and resources.
@@
To bridge this gap, we propose RTM (Requirements coverage-guided Test suite Minimization), a novel approach specifically designed for requirement-based testing. RTM aims to reduce test suite redundancy while ensuring full requirement coverage, making it particularly well-suited to meet the practical needs of our industrial partners, and critical systems that must demonstrate functional safety in general. RTM takes requirement-based test cases–composed of natural language test steps–as input and computes pairwise similarity between them. Specifically, we investigate three different preprocessing methods to normalize the test steps, followed by seven distinct text embedding techniques–both sentence-level and word-level–to transform the textual data into vector representations: Term Frequency–Inverse Document Frequency (TF-IDF), Universal Sentence Encoder (USE) (Cer et al., 2018), LongT5 (Guo et al., 2022), Amazon Titan Text Embedding V2 (AWS, 2024), Word2Vec (Mikolov et al., 2013), GloVe (Pennington et al., 2014), FastText (Bojanowski et al., 2017). We then measure the similarity between test cases using three different distance metrics: cosine similarity and Euclidean distance for sentence-level embeddings, and Word Mover’s Distance (WMD) (Kusner et al., 2015) for word-level embeddings. Based on the computed similarities, we employ a Genetic Algorithm (GA) (Luke, 2013) to search for the optimal test suite that minimizes similarity, adheres to a fixed budget, and ensures full requirement coverage. To help the GA efficiently find valid subsets, we devise three initialization strategies designed to generate an initial population consisting of subsets satisfying both coverage and budget constraints.
@@
RTM was evaluated on a real-world automotive dataset consisting of $736$ system test cases across seven test runs, along with $54$ requirements. The results demonstrate that RTM significantly outperforms two baseline techniques in terms of fault detection rate (FDR) under various minimization budgets. Furthermore, we explore the impact of the redundancy level of the test suites on FDR. Our findings show that: ① RTM consistently outperforms all baseline techniques across different redundancy levels; ② The level of redundancy has a strong effect on the achieved FDR, for all TSM techniques; ③ Especially at lower minimization budgets and for highly redundant test suites, there is still significant room for improvement between RTM and the theoretical maximum FDR.
@@
  • We validate RTM on automotive system requirements and test cases, and compare it with a number of baseline techniques for test suite minimization across varying minimization budgets, demonstrating its superior fault detection capability.
  • We further investigate the impact of redundancy levels of test suites on the effectiveness of TSM, offering new insights into the relationship between test suite redundancy and TSM fault detection performance.
  • To foster reproducibility and facilitate further research, we have open-sourced our replication package—including source code and experimental scripts—which is publicly available at https://anonymous.4open.science/r/RTM-5C98/. Industrial requirements and test cases are confidential.
@@ 2.1. Test Cases based on Automotive Requirements
Like other critical systems, modern automotive software systems are expected to comply with functional safety standards such as ISO 26262 (Siegl et al., 2010). To ensure compliance, system engineers derive and maintain extensive test suites based on system requirements, which are executed to validate system behavior against expected requirements. Typically, test cases are specified in natural language, following templates, and then manually converted into test code and implemented by experts.
The test cases in our dataset are collected from our industry partner and designed for testing automotive system requirements. As shown in Figure 1, these test cases are specified in natural language. Each test case consists of a sequence of diagnostic actions, structured into clearly labeled steps. These steps typically include setting up the preconditions for testing, creating the fault conditions, checking how the automotive system detects the issues through Diagnostic Trouble Codes (DTCs), and confirming the system properly clears the DTC. The detailed test actions involve reading and setting system variables, awaiting signal conditions, sending diagnostic requests, and checking system responses. To this end, the test cases exhibit the following characteristics:
@@
However, as the number of system requirements increases, the corresponding test suites can become excessively large, leading to dramatically increased testing costs. Moreover, as new requirements are added, test cases are often written without checking the existing ones. This can lead to functional overlap between test cases, thus increasing redundancy in the test suite.
@@ 2.2. Test Suite Minimization
Test suite minimization (TSM) aims to reduce the size of the test suite while ensuring important testing objectives such as requirement coverage, fault detection capability, and test case diversity (Harrold et al., 1993; von Ronne, 1999). Given that test suite diversity is reported to be positively correlated with fault detection capability (Hemmati et al., 2013), minimizing similarity among test cases can enhance the fault detection capability of the minimized test suite. When applied properly, TSM can significantly reduce testing costs and time while minimizing its impact on system quality or compliance.
Let =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ be the set of system requirements for an automotive system, and let $\mathcal{T}=\{t_{1},t_{2},\dots,t_{m}\}$ be the corresponding test suite. Each test case $t_{i}\in\mathcal{T}$ covers a subset of requirements denoted by $Cover(t_{i})\subseteq\mathcal{R}$. Note that in our dataset, each test case only covers one requirement, whereas each requirement can be covered by many test cases. This is because each test case is specifically designed to validate one scenario where the requirement is expected to hold. Moreover, covering all requirements is a must in automotive testing, to comply with functional safety standards (Siegl et al., 2010), which is also strictly required by our industry partners. Let $Sim(t_{i},t_{j}):\mathcal{T}\times\mathcal{T}\rightarrow[0,1]$ denote the similarity between test cases $t_{i}$ and $t_{j}$.
Given a test budget $\mathcal{M}\in(0,1]$, the goal is to select a subset $\mathcal{T}^{\prime}\subseteq\mathcal{T}$ such that:
  • Coverage: $\bigcup_{t\in\mathcal{T}^{\prime}}Cover(t)=\mathcal{R}$,
  • Budget: $|\mathcal{T}^{\prime}|=|\mathcal{T}|\times\mathcal{M}$,
  • Diversity: the total pairwise similarity among test cases in $\mathcal{T}^{\prime}$ is minimized.
Formally, this can be defined as the following optimization problem:
Given: test suite $\mathcal{T}$, requirement set $\mathcal{R}$, budget $\mathcal{M}\in(0,1]$,
coverage function $Cover(t)\subseteq\mathcal{R}$ for each $t\in\mathcal{T}$, and similarity function $Sim(t_{i},t_{j})$.
@@
This problem extends the classical set cover problem, which is known to be NP-hard, by introducing two additional constraints: a strict minimization budget and a diversity objective. In particular, the selected test subset must maintain full requirement coverage under a fixed budget while minimizing the overall similarity among test cases. As a result, solving this problem exactly is computationally intractable for large-scale industrial systems. Therefore, practical heuristics or approximation strategies are typically employed in automotive domains to achieve efficient and effective test suite minimization, especially when additional factors, such as fault detection capability, execution cost, or coverage redundancy, are also considered.
@@ 3. Approach
This section introduces RTM, a TSM approach that satisfies the three criteria outlined in Section 2.2. It targets test cases derived from system requirements, for the purpose of validation. In short, it leverages natural language processing (NLP) techniques and large language models (LLMs) as well as various distance functions to compute test case similarity and then employs evolutionary search to optimize minimization.
As illustrated in Figure 2, RTM first preprocesses test cases using three different strategies, then converts them into vector representations using a variety of text embedding techniques, including both sentence-level and word-level representations: TF-IDF, USE, LongT5, Amazon Titan Text Embedding V2, Word2Vec, GloVe, and FastText. Different distance functions are employed to compute the similarity values between test case embeddings, which are then used as guidance for the evolutionary search to find the optimal minimized test suite $\mathcal{T}^{\prime}$ with diverse test cases. The primary objective of RTM is to find the optimal $\mathcal{T}^{\prime}$ that maximizes fault detection capability by minimizing the similarity values between the test cases while reducing the testing cost, measured in terms of test suite size, as well as ensuring $100\%$ requirement coverage.
[…]
Figure 2. Overall Framework of RTM.
@@ 3.1. Test Case Preprocessing
Given that the test cases are expressed in natural language, different preprocessing strategies may affect the outcomes of the text embedding techniques. To investigate this, we implemented three levels of preprocessing:
@@ 3.2. Test Case Representation
After the preprocessing step, we transform the test cases into vector representations using various embedding techniques. We employ both word-level and sentence-level embedding techniques to capture semantic information from test cases. For word-level embeddings, we use Word2Vec, GloVe, and FastText, each of which converts every word in the test case to a $300-$dimensional vector.
For sentence-level embeddings, we employ TF-IDF, USE, LongT5, and Amazon Titan Embedding V2, each of which converts the entire test case content to a fixed-length vector representation that captures its overall semantic meaning, with the dimensionality varying depending on the embedding technique used.
@@ 3.3. Similarity Measure
To assess the similarity between test case pairs, we utilize different distance metrics based on the type of embedding used.
For word-level embeddings (i.e., Word2Vec, GloVe, and FastText), we use WMD (Kusner et al., 2015), which measures the minimum cumulative distance required to align the words in one test case with those in another. WMD leverages the underlying word embedding space and is particularly effective in capturing semantic similarity when the word order or exact word match is less important (Kusner et al., 2015).
For sentence-level embeddings (i.e., TF-IDF, USE, LongT5, and Amazon Titan Embedding V2), we applied both cosine similarity and Euclidean distance, which have been widely used in machine learning, information retrieval, and NLP tasks to measure similarity or dissimilarity between vectors (Gomaa and Fahmy, 2013; Huang et al., 2008; Li and Han, 2013; Salton and Buckley, 1988). However, they operate differently and are suited for different use cases: cosine similarity measures the angle between two vectors while ignoring the magnitude, whereas Euclidean distance measures the straight-line distance between two vectors. By applying different types of similarity metrics, we aim to explore how different vector representations and distance functions affect the ability of RTM to detect redundant test cases.
This step constructs a similarity matrix by computing the pairwise similarity between test cases. It takes the vector representations of test case pairs as input and calculates the similarity scores to quantify their similarity.

### v2

@@ Abstract.
As software systems evolve, test suites tend to grow in size and often contain redundant test cases. Such redundancy increases testing effort, time, and cost. Test suite minimization (TSM) aims to eliminate such redundancy while preserving key properties, such as requirement coverage and fault-detection capability. In this paper, we propose RTM (Requirement coverage-guided Test suite Minimization), a novel TSM approach designed for requirement-based testing (validation), which can effectively reduce test suite redundancy while ensuring full requirement coverage and a high fault detection rate (FDR) under a fixed minimization budget. Based on common practice in critical systems where functional safety is important, we assume that test cases are specified in natural language and traced to requirements before implementation. RTM utilizes text embedding technique to convert test cases into vector representations, on which a distance function is employed to compute similarity values between test case pairs. Guided by these similarity values, a Genetic Algorithm (GA) whose population is initialized using a coverage-preserving strategy is then employed to search for an optimal subset of diverse test cases that matches the budget. We investigate three preprocessing methods for test cases, seven different text embedding techniques, three distance functions, and three initialization strategies for the GA. We evaluate RTM on an industrial automotive system dataset comprising $736$ system test cases and $54$ requirements. Experimental results show that, while being scalable in terms of runtime, RTM outperforms all the baseline techniques in terms of FDR on most minimization budgets while maintaining full requirement coverage. Furthermore, we investigate the impact of test suite redundancy levels on the effectiveness of TSM, providing new insights into optimizing requirement-based test suites under practical constraints.
@@ 1. Introduction
Software testing is a critical activity throughout the software development life cycle, playing a key role in early defect detection and in ensuring requirements compliance (myers2011art; everett2007software). In particular, requirements-driven testing is conducted to validate systems before deployment (dustin2009implementing). This is especially critical for systems that must demonstrate functional safety, such as automotive systems. In such contexts, it is also common to specify test cases in natural language (following templates) to facilitate their analysis and traceability to requirements, before manually writing the corresponding test code.
@@
However, due to the increasing complexity of systems and their continuous change by large teams, test suites are hard to maintain, tend to grow, and are often redundant (i.e., some of the test cases are likely to detect the same faults), thus leading to waste of time and resources (rothermel2002empirical). Further, in practice, testers have limited resources and time to test new system updates. To address this, given a test budget, test suite minimization (TSM) is proposed to prune test cases that are most likely to be redundant, while satisfying coverage criteria and maximizing fault detection (harrold1993methodology; von1999test), thus significantly reducing testing cost and time while limiting the impact on system quality or compliance.
@@
While several approaches have been proposed for TSM (noemmer2019evaluation; miranda2017scope; coviello2018clustering; liu2011user; viggiato2022identifying; cruciani2019scalable; zhang2019uncertainty; hemmati2013achieving; pan2023atm; DBLP:journals/tse/PanGB24; marchetto2017combining; xia2021test; anwar2019hybrid; yoo2010using), most existing techniques rely on code and structural coverage criteria (e.g., statement or branch coverage). These techniques do not address the unique challenges of minimizing test suites driven by system requirements, with test cases often specified in natural language. Moreover, to the best of our knowledge, no existing solution can simultaneously maintain full requirement coverage while adhering to a fixed minimization budget, a common situation in industrial practice due to limited time and resources.
@@
To bridge this gap, we propose RTM (Requirements coverage-guided Test suite Minimization), a novel approach specifically designed for requirement-based testing. RTM aims to reduce test suite redundancy (in terms of fault detection) while ensuring full requirement coverage, making it particularly well-suited to the practical needs of our industrial partners and to critical systems that must demonstrate functional safety in general. RTM takes requirement-based test cases–composed of natural language test steps–as input and computes pairwise similarity between them. Specifically, we investigate three different preprocessing methods to normalize the test steps, followed by seven distinct text embedding techniques–both sentence-level and word-level–to transform the textual data into vector representations: Term Frequency–Inverse Document Frequency (TF-IDF), Universal Sentence Encoder (USE) (DBLP:journals/corr/abs-1803-11175), LongT5 (DBLP:conf/naacl/GuoAUONSY22), Amazon Titan Text Embedding V2 (aws2024titantv2), Word2Vec (DBLP:journals/corr/abs-1301-3781), GloVe (DBLP:conf/emnlp/PenningtonSM14), FastText (DBLP:journals/tacl/BojanowskiGJM17). We then measure the similarity between test cases using three different distance metrics: cosine similarity and Euclidean distance for sentence-level embeddings, and Word Mover’s Distance (WMD) (DBLP:conf/icml/KusnerSKW15) for word-level embeddings. Based on the computed similarities, we employ a Genetic Algorithm (GA) (Luke2013Metaheuristics) to search for the optimal test suite that minimizes similarity, adheres to a fixed budget, and ensures full requirement coverage. To help the GA efficiently find valid subsets, we devise three initialization strategies that generate an initial population of subsets satisfying both coverage and budget constraints.
@@
RTM was evaluated on an industrial automotive dataset consisting of $736$ system test cases across seven test runs, along with $54$ requirements. The results demonstrate that RTM outperforms the baselines in terms of fault detection rate (FDR) under various minimization budgets. Moreover, we explore the impact of test suite redundancy on FDR. Our findings show that: ① RTM consistently outperforms all baseline techniques in terms of FDR across different test suite redundancy levels; ② The level of redundancy has a strong effect on the achieved FDR, for all TSM techniques; ③ Especially at lower minimization budgets and for highly redundant test suites, there is still significant room for improvement between the theoretical maximum FDR and that of RTM. ④ The runtime of RTM scales approximately linearly with test suite size, with small (i.e., of no practical consequence) absolute differences compared to the baselines. Given its consistently higher FDR performance across various test suites with different redundancy levels, RTM is therefore a better choice in many practical contexts.
@@
  • We validate RTM on automotive system requirements and test cases, and compare it with several baseline techniques for test suite minimization across varying minimization budgets, demonstrating its superior fault detection capability.
  • We further investigate the impact of test suite redundancy levels on TSM effectiveness, offering new insights into the relationship between test suite redundancy and TSM fault-detection performance.
  • To foster reproducibility and facilitate further research, we have open-sourced our replication package, including source code, input data, and results for all research questions that allow the reproducibility of the experiments.
@@ 2.1. Test Cases based on Automotive Requirements
Like other critical systems, modern automotive software systems are expected to comply with functional safety standards such as ISO 26262 (siegl2010model). To ensure compliance, system engineers derive and maintain extensive test suites based on system requirements, which are executed to validate system behavior against expected requirements. Typically, test cases are specified in natural language, following templates, and then manually converted into test code and implemented by experts.
The test cases in our dataset were collected from our industry partner and designed to test automotive system requirements. As shown in Figure 1, these test cases are specified in natural language. Each test case consists of a sequence of diagnostic actions, structured into clearly labeled steps. These steps typically include setting up the preconditions for testing, creating the fault conditions, checking how the automotive system detects the issues through Diagnostic Trouble Codes (DTCs), and confirming that the system properly clears the DTC. The detailed test actions involve reading and setting system variables, awaiting signal conditions, sending diagnostic requests, and checking system responses. To this end, the test cases exhibit the following characteristics:
@@
However, as the number of system requirements increases, the corresponding test suites can become excessively large, resulting in dramatically higher testing costs. Moreover, as new requirements are added, test cases are often written without checking the existing ones. This can lead to functional overlap between test cases, thus increasing redundancy in the test suite.
@@ 2.2. Test Suite Minimization
Test suite minimization (TSM) aims to reduce the test suite size while ensuring key testing objectives, such as requirement coverage, fault-detection capability, and test-case diversity (harrold1993methodology; von1999test). Given that test suite diversity is reported to be positively correlated with fault detection capability (hemmati2013achieving), minimizing test case similarity can enhance the fault detection capability of the resulting test suite. When appropriately applied, TSM can significantly reduce testing costs and time while minimizing its impact on system quality or compliance.
Let $\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ be the set of system requirements for an automotive system, and let $\mathcal{T}=\{t_{1},t_{2},\dots,t_{m}\}$ be the corresponding test suite. Each test case $t_{i}\in\mathcal{T}$ covers a subset of requirements denoted by $Cover(t_{i})\subseteq\mathcal{R}$. Note that in our dataset, each test case only covers one requirement, whereas many test cases can cover each requirement. This is because each test case is specifically designed to validate a scenario in which the requirement is expected to hold. Therefore, in our context, guaranteeing full requirement coverage requires at least as many test cases as requirements in the test suites. Moreover, covering all requirements is a must in automotive testing, to comply with functional safety standards (siegl2010model), which are adhered to by our industry partner. Let $Sim(t_{i},t_{j}):\mathcal{T}\times\mathcal{T}\rightarrow[0,1]$ denote the similarity between test cases $t_{i}$ and $t_{j}$.
Specifically, TSM in our context can be defined as the following optimization problem:
Given: test suite $\mathcal{T}$, requirement set $\mathcal{R}$ with $r$ requirements, budget ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}\in(0,1]$, coverage function $Cover(t)\subseteq\mathcal{R}$ for each $t\in\mathcal{T}$, and similarity function $Sim(t_{i},t_{j})$.
@@
This problem extends the classical set cover problem, which is known to be NP-hard, by introducing two additional constraints: a strict minimization budget and a diversity objective. In particular, the selected test subset must maintain full requirement coverage within a fixed budget while minimizing overall test-case similarity. As a result, solving this problem exactly is computationally intractable for large-scale industrial systems. Therefore, practical heuristics or approximation strategies are typically employed in automotive domains to achieve efficient and effective test suite minimization, especially when additional factors, such as fault detection capability, execution cost, or coverage redundancy, are also considered.
@@ 3. Approach
This section introduces RTM, a TSM approach that satisfies the three criteria outlined in Section 2.2. It targets test cases derived from system requirements for validation. In short, it leverages a text embedding technique combined with a distance function to compute test case similarity, then employs evolutionary search to minimize the test suite. The primary objective of RTM is to find the optimal $\mathcal{T}^{\prime}$ that maximizes fault detection capability by minimizing the similarity values between the test cases while reducing the testing cost, measured in terms of test suite size, as well as ensuring $100\%$ requirement coverage.
Figure 2 outlines the main steps of RTM. We first describe how we preprocess test cases (Section 3.1) and then convert them into vector representations (Section 3.2). Then, we describe the distance function employed to calculate similarity values between test case embeddings (Section 3.3). Finally, we describe the search algorithm (GA) we employed to search for the near-optimal minimized test suite containing diverse test cases while satisfying all the constraints (Section 3.4).
Figure 2. Overall Framework of RTM.:
@@ 3.1. Test Case Preprocessing
Given that the test cases are expressed in natural language, different preprocessing strategies may affect the outcomes of the text embedding techniques. We investigated three different levels of preprocessing:
@@ 3.2. Test Case Representation
After the preprocessing step, we transform the preprocessed test cases into vector representations using a text embedding technique. We investigate both word-level and sentence-level embedding techniques to capture semantic information from test cases.
For word-level embeddings, we use Word2Vec, GloVe, and FastText, which maps each word in a test case to a $300$ dimensional numerical vector. Consequently, a test case with $k$ words is represented as a $k\times 300$ matrix. Note that we do not aggregate word vectors into a single sentence embedding. We retain the word-level embeddings, based on which the Word Mover’s Distance (WMD) is computed, as described in the following section.
For sentence-level embeddings, we employ TF-IDF, USE, LongT5, and Amazon Titan Embedding V2, each of which encodes the entire test case into a single fixed-length vector that captures the overall semantics of the entire test case content. The dimensionality of these vectors varies across embedding techniques. For example, TF-IDF produces sparse vectors whose dimensionality equals the vocabulary size, USE yields a $512$-dimensional vector, LongT5 produces a $768$-dimensional embedding, and Amazon Titan Embedding V2 returns a $1024$-dimensional vector. These differences do not affect our approach, as similarity, as described in the following section, is computed independently for each embedding technique without requiring dimensional alignment.
@@ 3.3. Similarity Measure
To assess the similarity between test case embeddings, we investigate different distance metrics depending on the embedding type.
For word-level embeddings (i.e., Word2Vec, GloVe, and FastText), we use WMD (DBLP:conf/icml/KusnerSKW15), which formulates the similarity between two test cases as an optimal transport problem. Specifically, WMD computes the minimum cumulative cost required to move the probability mass of all the words in one test case to those in another in the embedding space. The WMD is defined as:
(4) $\mathrm{WMD}(\mathbf{t}_{1},\mathbf{t}_{2})=\min_{T\geq 0}\sum_{i,j=1}^{n}T_{ij}\,d(i,j)$
where $n$ is the global vocabulary size, $T_{ij}$ denotes how much of the probability mass of word $i$ in test case $t_{1}$ moves to word $j$ in test case $t_{2}$, and $d(i,j)$ represents the Euclidean distance between word embeddings $i$ and $j$. If a word $i$ appears $f_{i}$ times in test case $t_{1}$, we denote the probability mass of word $i$ as $p_{1i}=\frac{f_{i}}{\sum_{j=1}^{n}f_{j}}$. The transport plan $T$ (a matrix) is subject to the constraints $\sum_{j}T_{ij}=p_{1i}$ and $\sum_{i}T_{ij}=p_{2j}$, $\forall i,j\in\{1,\dots,n\}$, which ensure test case $t_{1}$ is entirely transformed into test case $t_{2}$.
This optimization problem is equivalent to the Earth Mover’s Distance (monge1781memoire; wolsey1999integer; rubner1998metric) and can be solved via linear programming, with a computational complexity of approximately $O(n^{3}\log n)$ for $n$ unique words. Although the computation can be expensive when there are many words, WMD is well-suited to word-level embeddings because pairwise distances between the word embeddings approximate semantic dissimilarity, allowing the transport cost to reflect meaning differences between test cases (DBLP:conf/icml/KusnerSKW15).
In order to bound the WMD between 0 and $1$, and convert it from a distance function to a similarity measure, we normalize it as follows:
(5) $\operatorname{WMD}_{\text{norm}}(\mathbf{t}_{1},\mathbf{t}_{2})=\frac{1}{1+\operatorname{WMD}(\mathbf{t}_{1},\mathbf{t}_{2})}\in(0,1].$
For sentence-level embeddings (i.e., TF-IDF, USE, LongT5, and Amazon Titan Embedding V2), we investigate both cosine similarity and Euclidean distance, which have been widely used in machine learning, information retrieval, and NLP tasks to measure similarity or dissimilarity between vectors (gomaa2013survey; huang2008similarity; li2013distance; salton1988term). However, they operate differently and are suited to different use cases: cosine similarity measures the angle between two vectors, ignoring magnitude, whereas Euclidean distance measures the straight-line distance between two vectors. By applying different types of similarity metrics, we aim to explore how different vector representations and distance functions affect the ability of RTM to detect redundant test cases.
We normalize the cosine similarity and Euclidean distance as follows:
(6) $\operatorname{Cosine~Similarity}_{\text{norm}}(\mathbf{t}_{1},\mathbf{t}_{2})=1-\frac{\arccos{(cosine~similarity(\mathbf{t}_{1},\mathbf{t}_{2}))}}{\pi}=1-\frac{\arccos{\frac{\mathbf{T_{1}}\cdot\mathbf{T_{2}}}{\|\mathbf{T_{1}}\|\|\mathbf{T_{2}}\|}}}{\pi}\in[0,1].$
where $T_{1}$ and $T_{2}$ are the embeddings of test case $t_{1}$ and $t_{2}$, respectively.
(7) $\operatorname{Euclidean~Distance}_{\text{norm}}(\mathbf{t}_{1},\mathbf{t}_{2})=\frac{1}{1+Euclidean~Distance(\mathbf{t}_{1},\mathbf{t}_{2})}=\frac{1}{1+\sqrt{\sum_{i=1}^{n}(t_{1i}-t_{2i})^{2}}}\in(0,1].$
where $t_{1i}$ and $t_{2i}$ denote the $i^{th}$ element in the embeddings of test case $t_{1}$ and $t_{2}$, respectively. $n$ is the embedding dimension (i.e., total number of elements).
This step constructs a similarity matrix by computing the pairwise similarity between test cases. It takes the vector representations of test case pairs as input and calculates the similarity scores using one of the similarity measures to quantify their similarity.